OpenStudy (samigupta8):

the equation whose roots are the roots of x^7+3x^5+x^3-x^2+7x+2=0 with their signs changed is a.x^7+3x^5+x^3+x^2+7x+2=0 b.x^7+3x^5+x^3+7x-2=0 c.x^7+3x^5+x^3-x^2+7x-2=0

OpenStudy (samigupta8):

and in b option there is x^2 also after x^3

OpenStudy (anonymous):

B....

OpenStudy (samigupta8):

yes b option

OpenStudy (samigupta8):

@AllTehMaffs pls..hlp

OpenStudy (anonymous):

@Viyang

OpenStudy (anonymous):

if you have x to power of n just change the sign of the term x^(n-1) if it exists and change the sign of the integer at the end

OpenStudy (samigupta8):

then according to that we wud have b option as answer

OpenStudy (anonymous):

yes

OpenStudy (samigupta8):

y...

OpenStudy (anonymous):

I think it has to be a - for the last term either both x_0 terms are either positive or negative because it's positive, so if they *both change then that x_0 term *has to stay positive Right?

OpenStudy (samigupta8):

i did't get dat how u r getting a and even ans is a

OpenStudy (anonymous):

oohh, there are an odd number of roots. nm. My argument doesn't hold. Sorry :/

hartnn (hartnn):

to get roots with signed changed, just replace 'x' by '-x'

hartnn (hartnn):

sign of roots changed means -x will satisfy the resulting equation, so thats just f(-x)

hartnn (hartnn):

what do you get if you put x=-x in x^7+3x^5+x^3-x^2+7x+2=0

OpenStudy (samigupta8):

\[ -x^7-3x^5-x^3-x^2-7x+2=0\]

hartnn (hartnn):

yeah, now you can factor out the -1

OpenStudy (samigupta8):

x^7+3x^5+x^3+x^2+7x-2=0

hartnn (hartnn):

doubts ?

OpenStudy (samigupta8):

no